# Transforming Code Repositories into MCP Services

### v2

@@ A.3 Example MCP Tool Implementations Generated by Code2MCP
A.3.1 Tools Generated from the ESM Repository
@@
A.3.2 Tools Generated from the SymPy Repository
@@
MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression with respect to variable """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression with respect to variable """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
@@
MCP Tool for Fourier Transform @mcp.tool(name="fourier_transform") def fourier_transform(expression: str, time_var: str, freq_var: str): """ Calculate Fourier transform """ try: from calc.symbolic import sympify, symbols import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (-50, 50) omega_vals = np.linspace(-10, 10, 100) result = [] for omega_val in omega_vals: def f(t_val): return integrand(t_val) * np.exp(-1j * omega_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append(complex(integral_result)) return {"success": True, "result": [str(complex(r)) for r in result]} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Laplace Transform @mcp.tool(name="laplace_transform") def laplace_transform(expression: str, time_var: str, laplace_var: str): """ Calculate Laplace transform of a function. """ try: from calc.symbolic import sympify, symbols from calc.transforms import laplace_transform as sympy_laplace_transform import numpy as np from scipy import integrate expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) try: F = sympy_laplace_transform(expr, t, s, noconds=True) return {"success": True, "result": str(F)} except Exception: def integrand(t_val): return float(expr.subs(t, t_val)) t_range = (0, 100) s_vals = np.linspace(0.1, 10, 50) result = [] for s_val in s_vals: def f(t_val): return integrand(t_val) * np.exp(-s_val * t_val) integral_result, _ = scipy_integrate.quad(f, t_range[0], t_range[1], limit=100) result.append({"s": float(s_val), "F": str(complex(integral_result))}) return {"success": True, "result": result} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Z Transform @mcp.tool(name="z_transform") def z_transform(expression: str, time_var: str, z_var: str, limit: int = 100): """ Calculate Z transform """ try: def ser(x): if isinstance(x, Basic): return str(x) if isinstance(x, (list, tuple, set)): return [ser(i) for i in x] if isinstance(x, dict): return {k: ser(v) for k, v in x.items()} return x expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = Sum(expr * (z ** (-n)), (n, 0, $o$)).doit() result = simplify(result) return {"success": True, "result": ser(result)} except Exception: try: if expr.is_Pow and expr.base.is_symbol and expr.exp == n: a = expr.base result = z / (z - a) return {"success": True, "result": ser(result)} elif expr.is_Mul and len(expr.args) == 2: for arg in expr.args: if arg == n: other = expr / n if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z / (z - a)**2 return {"success": True, "result": ser(result)} elif expr.is_Mul and n**2 in expr.args: other = expr / n**2 if other.is_Pow and other.base.is_symbol and other.exp == n: a = other.base result = a * z * (z + a) / (z - a)**3 return {"success": True, "result": ser(result)} result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception: result = 0 for k in range(limit + 1): term = expr.subs(n, k) * (z ** (-k)) result += term return {"success": True, "result": ser(result)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
A.3.3 Tools Generated from the Foam-Agent Repository

### v3

@@ Appendix D Example MCP Tool Implementations Generated by Code2MCP
D.1 Tools Generated from the ESM Repository
@@
D.1.1 Tools Generated from the SymPy Repository
@@
MCP Tool for Differentiation @mcp.tool(name="differentiate") def differentiate(expr: str, variable: str): """ Calculate derivative of expression """ try: from sympy import sympify, symbols, diff, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = diff(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Integration @mcp.tool(name="integrate_expression") def integrate_expression(expr: str, variable: str): """ Calculate integral of expression """ try: from sympy import sympify, symbols, integrate, Basic def ser(x): if isinstance(x, Basic): return str(x) return x expr_sym = sympify(expr) var = symbols(variable) res = integrate(expr_sym, var) return {"success": True, "result": ser(res)} except Exception as e: return {"success": False, "result": None, "error": str(e)}
@@
MCP Tool for Fourier Transform from sympy import sympify, symbols, fourier_transform as sympy_fourier_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="fourier_transform") def fourier_transform_tool(expression: str, time_var: str = "t", freq_var: str = "w"): try: expr = sympify(expression) t = symbols(time_var) omega = symbols(freq_var) F = sympy_fourier_transform(expr, t, omega) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Laplace Transform from sympy import sympify, symbols, laplace_transform as sympy_laplace_transform, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="laplace_transform") def laplace_transform_tool(expression: str, time_var: str = "t", laplace_var: str = "s"): try: expr = sympify(expression) t = symbols(time_var) s = symbols(laplace_var) F, _, _ = sympy_laplace_transform(expr, t, s) return {"success": True, "result": _serialize(F), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
MCP Tool for Z Transform from sympy import sympify, symbols, summation, $o$, Basic def _serialize(obj): if isinstance(obj, Basic): return str(obj) if isinstance(obj, (list, tuple, set)): return [_serialize(x) for x in obj] if isinstance(obj, dict): return {k: _serialize(v) for k, v in obj.items()} return obj @mcp.tool(name="z_transform") def z_transform_tool(expression: str, time_var: str = "n", z_var: str = "z", limit: int = 100): try: expr = sympify(expression) n = symbols(time_var) z = symbols(z_var) try: result = summation(expr * z**(-n), (n, 0, $o$)) except Exception: result = summation(expr * z**(-n), (n, 0, limit)) return {"success": True, "result": _serialize(result), "error": None} except Exception as e: return {"success": False, "result": None, "error": str(e)}
D.1.2 Tools Generated from the Foam-Agent Repository